OpenStudy (anonymous):

Write the expression $\frac{4+3c}{7}+2$ as a single fraction.

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\huge \frac{4+3c}{7}+2\]?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

that was right

OpenStudy (anonymous):

you can clean up the numerator by distributing the two and combining like terms

OpenStudy (anonymous):

it says to make it to a single fraction

OpenStudy (anonymous):

ok lets make it in to a single fraction

OpenStudy (anonymous):

\[\frac{4+3c}{7}+2=\frac{4+3c+2\times 7}{7}\]